OpenStudy (anonymous):

(help me pleaseee. I'll give you a fan and a medal) What is the value of 2x^-2y^-2 for x = 3 and y=-2 a. 1/18 b. 72 c. 2(-6)^-4 d. 1/648

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

\[\frac{ 2 }{ x^2*y^2 }\]

OpenStudy (danjs):

the negative exponents, mean put them on the other side of the fraction and change the sign on the exponent \[\frac{ a^{-2} }{ 1 } = \frac{ 1 }{ a^2 }\]

OpenStudy (anonymous):

oki O_O wait what Im confused?

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

*confused again*

OpenStudy (danjs):

\[2*x^{-2}*y^{-2}\] those are negative exponents, you need to move them to the denominator of the fraction , then the exponents will become positive

OpenStudy (danjs):

\[\frac{ 2 }{ x^2*y^2 }\]

OpenStudy (anonymous):

so all I gotta do is put it in like that ^

OpenStudy (danjs):

right, then use the values they give x = 3 and y=-2 , in there and evaluate it

OpenStudy (danjs):

(-2)^2 = (-2)*(-2) = +4

OpenStudy (danjs):

\[\frac{ 2 }{ 3^2*(-2)^2 }\]

OpenStudy (anonymous):

I got 0.8888889

OpenStudy (danjs):

\[\frac{ 2 }{ 3^2*(-2)^2 } = \frac{ 2 }{ 9 * 4 } = \frac{ 2 }{ 36 } = \frac{ 1 }{ 18 }\]

OpenStudy (anonymous):

ohh okay. I see now

OpenStudy (danjs):

that look ok?

OpenStudy (anonymous):

yes it does :)